# Learning from Mistakes: Enhancing Harmful Meme Detection via Misjudgment Risk Patterns

### v1

@@ Abstract.
Internet memes have emerged as a popular multimodal medium, yet they are increasingly weaponized to convey harmful opinions through subtle rhetorical devices like irony and metaphor. Existing detection approaches, including MLLM-based techniques, struggle with these implicit expressions, leading to frequent misjudgments. This paper introduces PatMD, a novel approach that improves harmful meme detection by learning from and proactively mitigating these potential misjudgment risks. Our core idea is to move beyond superficial content-level matching and instead identify the underlying misjudgment risk patterns, proactively guiding the MLLMs to avoid known misjudgment pitfalls. We first construct a knowledge base where each meme is deconstructed into a misjudgment risk pattern explaining why it might be misjudged, either overlooking harmful undertones (false negative) or overinterpreting benign content (false positive). For a given target meme, PatMD retrieves relevant patterns and utilizes them to dynamically guide the MLLM’s reasoning. Experiments on a benchmark of 6,626 memes across 5 harmful detection tasks show that PatMD outperforms state-of-the-art baselines, achieving an average of 8.30% improvement in F1-score and 7.71% improvement in accuracy, demonstrating strong generalizability and improved detection capability of harmful memes.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; $X$, 2025; XXisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
With the widespread adoption of the Internet, memes have emerged as one of the most vibrant cultural phenomena of the digital age. These multimodal information units, which ingeniously combine images and text, leverage their humorous, satirical, and highly viral nature to proliferate across social media platforms. This phenomenon has profoundly reshaped the modes of information production and consumption. Beyond being a medium for entertainment and social interaction, memes have evolved into a potent tool for expressing viewpoints, fostering community cohesion, and even influencing public opinion (Lin et al., 2024). However, this powerful influence is a double-edged sword. Memes have also become a vessel for packaging and amplifying hate speech, reinforcing biases, and disseminating implicit discrimination and misinformation (Sharma et al., 2022), posing a severe threat to the health of the online platforms.
To safeguard online platforms from the pervasive negative impact of harmful memes, developing automated detection approaches has emerged as a critical area of research (Kiela et al., 2020). Initial efforts in this domain largely adopted a feature-fusion paradigm, first processing the image and text through independent encoders and then combining their representations to identify harmful content (Suryawanshi et al., 2020; Pramanick et al., 2021). Subsequently, the field shifted towards fine-tuning large pre-trained multimodal models, which jointly model visual and textual inputs to achieve superior cross-modal alignment (Lu et al., 2019; Chen et al., 2020). More recently, to address the need for deeper semantic reasoning, research has pivoted towards leveraging the capabilities of Multimodal Large Language Models (MLLMs), either by generating explicit reasoning chains (Hee et al., 2022; Lin et al., 2023) or by retrieving similar samples to augment the decision-making process (Liu et al., 2025).
@@
Figure 1. The motivation example of harmful meme detection with misjudgment risk patterns
While MLLMs demonstrate a strong capacity for content comprehension, they exhibit notable limitations in detecting harmful memes. They may correctly identify all elements within a meme, yet lack the conceptual reasoning capability necessary to navigate complex contextual cues—particularly when confronted with sarcasm, metaphor, or implicit stereotypes. This deficiency often leads to erroneous misjudgments. As shown in Figure 1, the MLLMs accurately identify the elements in the memes, such as humans and monkeys, but they cannot capture the underlying decision logic behind harmful intent, i.e., why & how the combination of elements causes racially hateful memes.
As previously mentioned, due to the prevalent use of rhetorical devices such as irony and analogy in harmful memes, MLLMs frequently commit misjudgments in specific, systematic ways. Our core insight is to preemptively address this issue by abstracting these failure modes into misjudgment risk patterns. These patterns serve as explicit guidance, cautioning the MLLM to guard against these common pitfalls during its final decision-making. As such, our work presents a paradigm shift: moving from a purely content-centric detection to a model decision-centric one. By focusing not only on whether a meme is harmful but also on how the model reasons about it and why that reasoning may fail, we directly mitigate characteristic cognitive and interpretive limitations of MLLMs in meme analysis. As illustrated in Figure 1, we derive the misjudgment risk pattern from similar memes, indicating that the racially hateful information comes from the incorrect human and non-human comparison and causes a deliberate dehumanization. With the help of this pattern, MLLMs go beyond the separated human and animal elements and capture the underlying decision logic from human-animal relations. The misjudgment is mitigated, and MLLMs successfully identify the racially hateful memes.
@@
In this paper, we propose PatMD, an automated approach that leverages misjudgment risk patterns to improve harmful meme detection. PatMD constructs a knowledge base by diagnostically decomposing memes to archive past MLLM misjudgments. For any new meme, it retrieves these misjudgment risk patterns and injects them the MLLM’s reasoning process, proactively steer the analysis away from known pitfalls. PatMD consists of three stages: In misjudgment risk pattern elicitation stage, historical memes are mapped into a hierarchical harm tree, capturing multi-granular harmful concepts, from which corresponding misjudgment risk patterns are derived. Then risk-aware pattern retrieval stage identifies semantically similar memes based on structural similarity within the hierarchical harm tree, and fetches their associated risk patterns. In pattern-augmented reasoning for the detection stage, the retrieved patterns are integrated into a structured prompt, guiding the MLLMs to avoid known misjudgment pitfalls.
@@
To evaluate the performance of PatMD, we conduct extensive experiments on a benchmark with 6,626 memes spanning 5 harmful detection tasks. The results demonstrate that PatMD have substantial and consistent improvements against four specialized baselines and six general-purpose MLLMs, with an average improvement of 8.30% in F1-score and 7.71% in accuracy. The ablation study reveals the necessity of each component in our approach design. Further analysis reveals that the strength of PatMD lies in its ability to steer MLLMs toward proactive misjudgment prevention by focusing on the underlying logic of harm manifestation, rather than relying on superficial content analysis. This misjudgment-aware reasoning paradigm enables the model to perform more accurate and nuanced judgment, effectively reducing prevalent misclassification errors.
@@
  • We propose PatMD, a novel approach that improves harmful meme detection by learning from potential misjudgment risk patterns of historical memes. This presents a paradigm shift for harmful meme detection: moving from a purely content-centric detection to a model decision-centric one, i.e., proactively steering the MLLM’s reasoning away from known misjudgment pitfalls.
  • We conduct comprehensive experiments on a benchmark covering five harmful detection tasks, comparing with four specialized baselines and six general-purpose MLLMs. The results demonstrate its effectiveness, achieving an average of 8.30% improvement in F1-score and 7.71% improvement in accuracy.
  • Our code and dataset are available  [^1] https://anonymous.4open.science/r/PatMD-61BF for replication.
@@ 2. Related Works
Harmful meme detection has developed into an important direction in multimodal research and has continued to advance with the introduction of large-scale benchmark datasets (Cai et al., 2019; Kiela et al., 2020; Fersini et al., 2022; Lin et al., 2024). Early studies mainly focused on feature fusion, typically adopting a two-stream architecture that extracts visual and textual representations through independent encoders and then leverages attention mechanisms for cross-modal interaction and alignment (Suryawanshi et al., 2020; Pramanick et al., 2021; Lee et al., 2021). Subsequently, research shifted towards shared Transformer-based intermediate fusion and fine-tuning of pretrained multimodal models, both of which jointly model image regions and text sequences to enhance cross-modal alignment and discrimination (Kiela et al., 2019; Lu et al., 2019; Li et al., 2019; Chen et al., 2020). With the introduction of the Hateful Memes Challenge (Kiela et al., 2020), a variety of ensemble-based and contrastive reweighting approaches emerged (Muennighoff, 2020; Sandulescu, 2020; Lippe et al., 2020), further improving their generalizability. Other researchers explore prompt-based methods. Some work in this area (Cao et al., 2022; Ji et al., 2023; Cao et al., 2023; Ji et al., 2024) focused on concatenating the meme’s text with extracted image captions to fine-tune pre-trained language models(Liu et al., 2019; Chung et al., 2024) for the detection task. Additionally, recent work (Lin et al., 2025) has employed an interactive, multi-agent framework for deeper visual understanding. This framework simulates a discussion where agents query a vision expert to generate a synthesized analysis.
Figure 2. The overview of PatMD:
@@
Large Language Models (LLMs) have demonstrated exceptional capabilities in complex reasoning (Brown et al., 2020; Chowdhery et al., 2023), such as generating intermediate reasoning steps via Chain-of-Thought (CoT) prompting to enhance judgment accuracy (Kojima et al., 2022; Wei et al., 2022; Zhang et al., 2022). Inspired by this, research in interpretable harmful meme detection has begun to leverage LLMs to generate explicit chains of reasoning to explain their rationale. These methods, such as constructing LLM debate frameworks (Hee et al., 2022) or distilling LLM’s reasoning knowledge into smaller models (Lin et al., 2023). To address challenges in low-resource generalization, researchers have explored few-shot enhanced methods (Cao et al., 2024; Huang et al., 2024). Inspired by this direction, Liu et al. introduced MIND (Liu et al., 2025), a framework for zero-shot harmful meme detection that integrates similar example retrieval and multi-agent collaboration. This aligns with the broader paradigm of Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Guu et al., 2020), which enriches an LLM’s input with external knowledge and has proven highly effective in knowledge-intensive tasks (Ram et al., 2023; Jiang et al., 2023; Asai et al., 2024; Wei et al., 2024). While the aforementioned approaches improve performance by enhancing content understanding or providing content examples, they remain fundamentally content-centric. When adversaries design novel memes, these methods may struggle to find relevant content for reference. Inspired by these lines of research, we propose a paradigm shift. Instead of focusing on ”what the content is,” our method focuses on ”why the model might misjudge.” We introduce an approach based on misjudgment risk patterns and integrate it into the LLM’s decision-making process. This model-decision-centric perspective allows it to identify and guard against specific reasoning paths that lead to errors.
@@ 3. Overview of PatMD
Our approach, PatMD, is designed to enhance the harmful meme detection capabilities of MLLMs. We shift the paradigm from reliance on the MLLM’s implicit knowledge to proactively guide it with explicit misjudgment risk patterns derived from historical memes. These patterns capture the specific manner in which the MLLM may misinterpret the multimodal content. During inference, relevant patterns are used to steer the MLLM’s reasoning process, proactively alerting it to potential pitfalls and thereby enabling more robust judgments. The implementation of PatMD is structured into three main stages, as demonstrated in Figure 2.
@@ 3.1.1. Hierarchical Harm Deconstruction
Hierarchical Harm Tree Structure. The tree incorporates a root node $RNode$ and other tree nodes $TNode$. If the tree node is at the bottom level, it is considered a leaf node $LNodes$. The tree can be formalized as follows:
@@
(2) $\mathcal{T}_{\text{H}}^{(i)}=\text{MLLM}_{\text{gen}}(m_{i},\text{Template}(\mathcal{T}_{\text{H}}))$
where for a given meme $m_{i}$, the generation function $\text{MLLM}_{\text{gen}}$ fills the nodes of the tree template $\text{Template}(\mathcal{T}_{\text{H}})$ to produce the final instantiated tree $\mathcal{T}_{\text{H}}^{(i)}$ (See Appendix C Figure 7 for the detailed prompt templates). This process yields a structural harm instance in the hierarchical harm tree. For instance, as shown in Figure 3, the MLLM would identify the ‘Macro Category’ ($L_{1}$) as “Race”, the ‘Core Subject’ ($L_{2}$) as “African descent”, and so on.
Figure 3. Example of constructing the hierarchical harm tree and corresponding misjudgment risk pattern of the meme.
@@ 3.1.2. Causal Misjudgment Attribution
For benign samples, PatMD infers the root cause of its false positive risk based on the structural harm information. It analyzes why the combination of its theme, entities, and expressive techniques might be misinterpreted or over-interpreted as harmful.
For harmful samples, PatMD reasons in reverse, determining how the meme’s compositional elements are subtly employed to encode or conceal its false negative risk—that is, its underlying harmful intent.
@@
After the above process, each sample $m_{i}=(\text{img}_{i},\text{txt}_{i})$ is endowed with a structural harm instance $\mathcal{T}_{i}$ within the hierarchical harm tree and a misjudgment risk pattern $\mathcal{P}_{i}$. To make these information retrievable, we also employ a pre-trained multimodal encoder $E$ (Günther et al., 2025) to generate embedding vectors for its image and text, $v^{\text{img}}_{i}=E(\text{img}_{i})$ and $v^{\text{txt}}_{i}=E(\text{txt}_{i})$, respectively. This stage ultimately produces a knowledge base where each meme is encoded as a tuple $(v^{\text{img}}_{i},v^{\text{txt}}_{i},\mathcal{T}_{i},\mathcal{P}_{i},\mathcal{G}_{i})$, recording its multimodal embeddings, structural harm instance, misjudgment risk pattern, and ground truth label (benign/harmful) $\mathcal{G}_{i}$.

### v3

# Fall into a Pit, Gain in a Wit: Cognitive-Guided Harmful Meme Detection via Misjudgment Risk Pattern Retrieval

@@ Abstract.
Internet memes have emerged as a popular multimodal medium, yet they are increasingly weaponized to convey harmful opinions through subtle rhetorical devices like irony and metaphor. Existing detection approaches, including Multimodal Large Language Model (MLLM)-based techniques, struggle with these implicit expressions, leading to frequent misjudgments. This paper introduces PatMD, a novel approach that detects harmful memes by learning from and proactively mitigating these potential misjudgment risks. Our core idea is to move beyond superficial content-level matching and instead identify the underlying misjudgment risk patterns, proactively guiding the MLLMs to avoid known misjudgment pitfalls. We first construct a knowledge base where each meme is deconstructed into a misjudgment risk pattern explaining why it might be misjudged, either overlooking harmful undertones (false negative) or overinterpreting benign content (false positive). For a given target meme, PatMD retrieves relevant patterns and utilizes them to dynamically guide the MLLM’s reasoning. Experiments on a benchmark of 6,626 memes across 5 harmful detection tasks show that PatMD outperforms state-of-the-art baselines, achieving an average of 8.30% improvement in F1-score and 7.71% improvement in accuracy, while exhibiting consistent robustness on unseen and adversarial memes.
@@
Multimodal Large Language Model, Harmful Meme Detection.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Computer visionccs: Computing methodologies Natural language processing
@@ 1. Introduction
Nowadays, Internet memes have emerged as one of the most vibrant cultural phenomena of the digital age. They combine images and text, leveraging their humorous and satirical content to express viewpoints, fostering community cohesion, and even influencing public opinion (Lin et al., 2024). However, this powerful influence is a double-edged sword, because the widespread nature of memes makes them a vehicle for amplifying hateful speech, reinforcing biases, and disseminating discrimination (Sharma et al., 2022), posing a threat to online platforms.
@@
Figure 1. The motivation example of harmful meme detection with misjudgment risk patterns.
To detect harmful memes and safeguard online platforms, previous research has evolved from feature-fusion methods (Suryawanshi et al., 2020; Pramanick et al., 2021; Lu et al., 2019; Chen et al., 2020) to MLLM-based approaches (Hee et al., 2022; Lin et al., 2023, 2025), including recent Retrieval-Augmented Generation (RAG) paradigms (Huang et al., 2024; Liu et al., 2025) that reference historical memes to assist decision-making. However, these advanced methods remain fundamentally content-centric. They primarily rely on superficial visual and textual similarity, which makes them struggle when novel memes employ entirely different templates to convey similar implicit rhetoric (e.g., sarcasm or metaphor) (Kiela et al., 2020; Lin et al., 2024). More importantly, these prior works fail to formalize the MLLM’s inherent reasoning pitfalls into generalized and reusable misjudgment risk pattern that explicitly explain the underlying reasons for potential misjudgments.
Figure 1 illustrates the contribution of cognitive blind spot. The upper part of MLLM’s response shows that while standard MLLM can identify isolated elements (e.g., a human and a monkey), they fail to capture why and how their juxtaposition constitutes a racially hateful analogy. In contrast, our approach introduces a generalized misjudgment risk pattern that incorporates the human cognitive and decision-making process in analyzing harmful memes, proactively warning the model of derogatory human-animal comparisons intended to deliberate dehumanization. Equipped with this structured thought, the MLLM transcends isolated visual features to comprehend the implicit rhetorical relations, successfully mitigating the misjudgment and identifying the harmful meme.
Based on this thought of misjudgment risk pattern, our work presents a paradigm shift from purely content-centric detection to a cognitive decision-centric approach. Our core idea is to preemptively abstract systematic failure modes into misjudgment risk patterns. By explicitly cautioning the MLLM against common cognitive pitfalls (e.g., over-interpreting irony or missing analogies) during its reasoning, this approach naturally fosters a more transparent decision logic. However, transitioning to this paradigm entails two key challenges: (1) Historical misjudgments are scattered and unstructured, necessitating a systematic framework to categorize and archive these cognitive blind spots; (2) There lacks an effective mechanism to extract these implicit reasoning patterns from past errors and accurately retrieve them for novel memes.
In this paper, we propose PatMD, an automated approach that leverages misjudgment risk Patterns to improve harmful Meme Detection. Inspired by the concept “Fall into a Pit, Gain in a Wit”, where “Pit” refers to the inherent misjudgment traps that models often encounter when processing memes, and “Wit” represents the derived cognitive patterns used to steer the MLLM away from these pitfalls. PatMD first constructs a knowledge base by diagnostically decomposing memes to archive past MLLM misjudgments. For any target meme, it retrieves these misjudgment risk patterns and injects them into the MLLM’s reasoning process, steering the unknown analysis away from known pitfalls:
@@
In misjudgment risk pattern elicitation stage, historical memes are mapped into a hierarchical harm tree, capturing multi-granular harmful concepts, from which corresponding misjudgment risk patterns are derived. Then risk-aware pattern retrieval stage identifies semantically similar memes based on structural similarity within the hierarchical harm tree, and fetches their associated risk patterns. In pattern-augmented reasoning for the detection stage, the retrieved patterns are integrated into a structured prompt, guiding the MLLM to avoid known misjudgment pitfalls.
@@
To evaluate the performance of PatMD, we conduct extensive experiments on a benchmark with 6,626 memes spanning 5 harmful detection tasks. The results demonstrate that PatMD has substantial and consistent improvements against four specialized baselines and six general-purpose MLLMs, with an average improvement of 8.30% in F1-score and 7.71% in accuracy. The ablation study reveals the necessity of each component in our approach design. Moreover, PatMD demonstrates exceptional generality and robustness, effectively adapting to unseen domains and resisting adversarial visual perturbations. Further analysis reveals that the strength of PatMD lies in its ability to steer MLLM toward proactive misjudgment prevention by focusing on the underlying logic of harm manifestation, rather than relying on superficial content analysis. This misjudgment-aware reasoning paradigm enables the model to perform more accurate and nuanced judgments, effectively reducing prevalent misclassification errors.
@@
  • We propose PatMD, a cognitive-guided approach that improves harmful meme detection by learning from misjudgment risk patterns of historical memes, which proactively steers the MLLM’s reasoning away from misjudgment pitfalls.
  • We conduct experiments on a benchmark covering five harmful detection tasks. The results demonstrate that PatMD outperforms baselines on F1-score with high consistent robustness against unseen and adversarial memes.
@@ 2. Related Works
Harmful meme detection has developed into an important direction in multimodal research and has continued to advance with the introduction of large-scale benchmark datasets (Cai et al., 2019; Kiela et al., 2020; Fersini et al., 2022; Lin et al., 2024). Early studies mainly focused on feature fusion, typically adopting a two-stream architecture that extracts visual and textual representations through independent encoders and then leverages attention mechanisms for cross-modal interaction and alignment (Suryawanshi et al., 2020; Pramanick et al., 2021; Lee et al., 2021). Subsequently, research shifted towards shared Transformer-based intermediate fusion and fine-tuning of pretrained multimodal models, both of which jointly model image regions and text sequences to enhance cross-modal alignment and discrimination (Kiela et al., 2019; Lu et al., 2019; Li et al., 2019; Chen et al., 2020). With the introduction of the Hateful Memes Challenge (Kiela et al., 2020), a variety of ensemble-based and contrastive reweighting approaches emerged (Muennighoff, 2020; Sandulescu, 2020; Lippe et al., 2020), further improving their generalizability. Other researchers explore prompt-based methods. Some work in this area (Cao et al., 2022; Ji et al., 2023; Cao et al., 2023; Ji et al., 2024) focused on concatenating the meme’s text with extracted image captions to fine-tune pre-trained language models(Liu et al., 2019; Chung et al., 2024) for the detection task.
A flow diagram showing the modules of our proposed method, including construct, retrieve, and inference.
Figure 2. The overview of PatMD.
@@
Recently, MLLMs have demonstrated exceptional capabilities in complex reasoning (Brown et al., 2020; Chowdhery et al., 2023), such as generating intermediate reasoning steps via Chain-of-Thought prompting to enhance judgment accuracy (Kojima et al., 2022; Wei et al., 2022; Zhang et al., 2022). Inspired by this, research in interpretable harmful meme detection has begun to leverage LLMs to generate explicit chains of reasoning to explain their rationale. These methods, such as constructing LLM debate frameworks (Hee et al., 2022) or distilling LLM’s reasoning knowledge into smaller models (Lin et al., 2023). Building upon this interactive paradigm, recent work (Lin et al., 2025) has employed a multi-agent framework that simulates a discussion where agents query a vision expert for deeper visual understanding. To address challenges in low-resource generalization, researchers have explored few-shot enhanced methods (Cao et al., 2024; Huang et al., 2024). Inspired by this direction, Liu et al. introduced MIND (Liu et al., 2025), a framework for zero-shot harmful meme detection that integrates similar example retrieval and multi-agent collaboration. This aligns with the broader paradigm of RAG (Lewis et al., 2020; Guu et al., 2020), which enriches an LLM’s input with external knowledge and has proven highly effective in knowledge-intensive tasks (Ram et al., 2023; Jiang et al., 2023; Asai et al., 2024; Wei et al., 2024).
While the aforementioned approaches improve performance by enhancing content understanding or providing content examples, they remain fundamentally content-centric. When malicious users craft novel and implicit memes, these methods may struggle to find relevant content for reference. Inspired by these lines of research, we propose a paradigm shift. Instead of focusing on “what the content is,” our method focuses on “why the model might misjudge.” We introduce an approach based on misjudgment risk patterns and integrate it into the LLM’s decision-making process. This cognitive decision-centric perspective allows it to identify and guard against specific reasoning paths that lead to errors.
@@ 3. Overview of PatMD
Figure 2 illustrates the structure of PatMD, which formalizes the flow “Memes Base $\rightarrow$ Pattern Elicitation $\rightarrow$ Pattern Retrieval $\rightarrow$ Detection”, to detect harmful memes by integrating MLLM’s capabilities within a RAG paradigm. We shift the paradigm from reliance on the MLLM’s implicit knowledge to proactively guide it with explicit misjudgment risk patterns derived from historical memes. These patterns capture the specific manner in which the MLLM may misinterpret the multimodal content. During inference, relevant patterns are used to steer the MLLM’s reasoning process, proactively alerting it to potential pitfalls and thereby enabling more robust judgments. The implementation of PatMD is structured into three main stages:
@@ 3.1.1. Hierarchical Harm Deconstruction
Hierarchical Harm Tree Structure Definition. The tree incorporates a root node $RNode$ and other tree nodes $TNode$. If the tree node is at the bottom level, it is considered a leaf node $LNodes$. The tree can be formalized as follows:
@@
Specifically, for a given meme $m_{i}$, the generation function $\text{MLLM}_{\text{gen}}$ fills the nodes of the tree template $\text{Template}(\mathcal{T}_{\text{H}})$ to produce the final instantiated tree, formulated as $\mathcal{T}_{\text{H}}^{(i)}=\text{MLLM}_{\text{gen}}(m_{i},\text{Template}(\mathcal{T}_{\text{H}}))$ (See Appendix C Figure 9 for the detailed prompt templates). This process yields a structural harm instance in the hierarchical harm tree. For instance, as shown in Figure 3, the MLLM would identify the ‘Macro Category’ ($L_{1}$) as “Disability”, the ‘Core Subject’ ($L_{2}$) as “Child in a wheelchair”, and so on.
Figure 3. The example of PatMD hierarchical harm tree and structural matching versus content matching.
@@ 3.1.2. Causal Misjudgment Attribution
  • For benign samples, PatMD infers the root cause of its false positive risk based on the structural harm information. It analyzes why the combination of its theme, entities, and expressive techniques is misinterpreted or over-interpreted as harmful.
  • For harmful samples, PatMD reasons in reverse, determining how the meme’s compositional elements are subtly employed to encode or conceal its false negative risk, which is its underlying harmful intent.
@@
After the above process, each sample $m_{i}=(\text{img}_{i},\text{txt}_{i})$ is endowed with a structural harm instance $\mathcal{T}_{i}$ within the hierarchical harm tree and a misjudgment risk pattern $\mathcal{P}_{i}$. To make these information retrievable, we also employ a pre-trained multimodal encoder $E$ (Günther et al., 2025) to generate embedding vectors for its image and text, $v^{\text{img}}_{i}=E(\text{img}_{i})$ and $v^{\text{txt}}_{i}=E(\text{txt}_{i})$, respectively.
This stage ultimately produces a knowledge base where each meme is encoded as a tuple $(v^{\text{img}}_{i},v^{\text{txt}}_{i},\mathcal{T}_{i},\mathcal{P}_{i},\mathcal{G}_{i})$, recording its multimodal embeddings, structural harm instance, misjudgment risk pattern, and ground truth label (benign/harmful) $\mathcal{G}_{i}$.